OpenStudy (anonymous):

The scale of a map is 1 cm : 25 km. Find the actual distance for a map distance of 6.2 cm.

OpenStudy (hoblos):

= 6.2 * 25 = 155km

OpenStudy (anonymous):

155 km....

OpenStudy (anonymous):

6.2 x 25 = (6 x 25) + (25 / 5) = 150 + 5 = 155 km